A rectangular garden is 5 ft longer than it is wide. Its area is 6000 ft{}^2. What are its dimensions?
Answer by jorel1380(3719)

n(n+5)=6000
n2+5n-6000=0
(x+80)(x-75)=0
x=-80 or 75
Throwing out the negative answer, we get the dimensions of the garden to be 75' by 80'..